What is the PH of a 100 mM solution of sulfuric acid (H2SO4)?

Ka1=Large

Ka2=1.2x10^-2

A. 2.02

B. -0.96

C. 1.00

D. 1.9

E. 0.96
Answer #1

[H2504 ] = 100mm = 0.10 .. (1000 mm = 1) completely means during H₂SO4 dimoclates first disociation Kai is large it in Hi andx= -0.112 + (0.112)2 - 4X 1 X (-0.0012) x = 9,85X10-3M [ht] =(0.10 + x) = (0.10+ 9.85 × 10-3) M = 0.10985 M = 0.96 pH = -log[